OpenStudy (anonymous):

write an equation to the line perpendicular to the line given by 2x+3y=7 and having the same y-intercept as the line given by 5x+2y=10.

OpenStudy (aroub):

3y=-2x+7 y= -2/3x+7/3 m= 3/2 ---------------------- 2y=-5x+10 y=-5/2x+5 b=5 ------------- the equation is y=3/2x+5

OpenStudy (anonymous):

yeah since youd take the opposite reciprocal of y =-2/3x and take the y int 5 from y=5-5/2x